### v2

# ”SNe Ia twins” in the Hubble flow, and the determination of H0

Abstract
We have applied our approach of using ”SNe Ia twins”in the Hubble flow to obtain distances to SNe Ia at z 0.015 and derive H0. Our results, taking a single step between the low z domain and the Hubble flow, validate the three rung classical method. We find, however, that the full compilation of distances, both in Pantheon+ and in the Carnegie-Chicago Hubble Program (CCHP), contain some inaccurate values in the colors due to an undersestimate of reddening by dust, or due to the adoption of not well–defined light curve declines. This produces odd individual values for H0 from single SNe Ia. On the average, those erroneous estimates do not affect the mean value of H0, which is characterized by the bulk of well–modeled SNe Ia. Our sample of carefully addressed SNe Ia in the Hubble flow contains a dozen supernovae, for which the distances are determined with high accuracy. Three of these SNe Ia are of the Broad Line subtype and can be compared with SN 1989B in M66, a host galaxy with a unique convergence of the Cepheid distance determination and the Tip of the Red Giant Branch stars (TRGB) determination by the CCHP group. They point to a weighted average of H0 73.556 2.084 (stat) km s-1 Mpc-1. There is as well a very good agreement on the distances to NGC 7250 and NGC 5643 between those derived with Cepheids by SH0ES and those derived with the use of J-Asymptotic Giant Branch stars (JAGB stars) by the CCHP, which makes them very good anchors. The sample of 12 SNe Ia gives a value of H0 72.833 1.306(stat) 1.151 (sys) km s-1 Mpc-1, when anchored in Cepheids, and of H0 72.388 1.272 (stat) 1.015 (sys) km s-1 Mpc-1, when anchored in JAGBs by the CCHP. We take a mean of the two values of H0 as derived by the Cepheids and by JAGB (from the CCHP) and obtain H0 72.610 1.289(stat) 1.085 (sys) km s-1 Mpc-1. Our findings confirm that the Hubble tension is real.
y
1 Introduction
Type Ia supernovae (SNe Ia) at high led to the discovery of the acceleration of the Universe (Riess et al. 1998; Perlmutter et al. 1999) and dark energy. Large compilations of SNe Ia have recently suggested a tentantive evolution in time of such unkown major component of the energy density of the Universe (Rubin et al. 2023; DESyr5, DESI111SH0ES stands for ”Supernova H0 for the Equation of State of Dark Energy”. DESyr5 for ”Dark Energy Survey year 5”, DESI for ”Dark Energy Spectroscopic Instrument”.).
Apart from this use in cosmology, a puzzle emerged some years ago and still requires an explanation: the expansion rate of the Universe today as obtained with SNe Ia calibrated with Cepheids gives a value of H0 higher than that measured from the Cosmic Microwave Background (CMB). The value of H0 derived from the CMB is 67.40.5 km s-1 Mpc-1 (Planck Collaboration 2020) and the latest SH0ES value of H0 = 73.290.9 km s-1 Mpc-1 (Murakami et al et. 2023) have now a discrepancy at the 5.7 level or even at the 6 level (Riess et al. 2025). The SH0ES group has now addressed three additional methods: their own Tip of the Red Giant Branch stars (TRGB) measurements, to be compared with that of the Carnegie Chicago Supernova Project (CCHP), their own J-Asymptotic Giant Branch stars (JAGB) and the use of Mira stars to compare with Cepheids, in conjunction with SNe Ia in the Hubble flow from the Pantheon+ SN catalog (Scolnic et al. 2022). They obtain values of 72.1-73.3 1.8 km s-1 Mpc-1 (depending on methodology) (Li et al. 2025).
Such difference in H0, when comparing with the H0 provided by the CMB, if confirmed, would stand as a challenge to the CDM model (Di Valentino et al. 2025a,b). A number of suggestions have been made, including the need of new physics in the early Universe to make the Planck value compatible with that derived by methods involving low- astrophysical distance indicators such as Cepheids (see Di Valentino et al. 2021a,b, 2025a,b; Kamionkowski & Riess 2023; Schöneberg et al. 2021; Mörtsell & Dhawan 2018, Poulin et al. 2025, for reviews).
Nowadays, several other methods are giving values of H0 around 74 km s-1 Mpc-1 : the sample of distances obtained by the Surface Brightness Fluctuations (SBF) method, dominated by early-type galaxies as required by the method, calibrated with TRGB, gives rise to an H0 73.8 0.7 (stat) 2.3 (syst) km s-1 Mpc-1 (Jensen et al. 2025). Garnavich et al. (2023) had used as well the SBF method and found H0 74.6 0.9 (stat) 2.7 (syst) km s-1 Mpc-1. The Expanding Photosphere method, using SNe II, gives 74.9 1.9 (stat) km s-1 Mpc-1 (Vogl et al. 2024).
While new methods are now providing the above quoted values for H0, the debate on obtaining H0 using Cepheids, TRGB or JAGB stars distances to calibrate Type Ia supernovae (SN Ia) persists. The SNe Ia calibration depends on three rungs: Rung 1 is the calibration of TRGB/Cepheids/JAGB stars with local anchors using geometrical methods, rung 2 refers to the calibration of the distance to SNe Ia host galaxies with TRGB/Cepheids/JAGB stars, and rung 3 is the overall calibration of SNe Ia in the Hubble flow, which is done in different ways but usually involves the derivation of a fiducial value for the absolute magnitude of SNe Ia and the final dependence of each particular SN Ia magnitude from such absolute magnitude and from factors such as rate of decline (hereafter called stretch) of the light curve, color of the SN Ia at maximum and mass of its host galaxy. The two collaborations are making this step in different ways: Pantheon+ uses the SALT2 relation (Brout et al. 2022) while the CCHP uses a polynomial derivation (Freedman et al. 2019).
Given the discrepancies between the two groups on H0, we decided to introduce a new approach (Ruiz–Lapuente & Gonzalez Hernández 2024; hereafter RLGH24). We address the Hubble tension issue with a method which uses the spectral evolution of SNe Ia that have not only similar stretch but identical spectral features along their lifetimes, i.e. ”SNe Ia twins for life”. The precision of the distance provided by those twins should be higher than for SNe Ia with just similar stretch. The total reddening could also be obtained.
The selection of twins is made of SNe Ia with a similar stretch, being then of similar luminosities, but in addition the ”twinness factor” can make more precise the relative distance estimate, with a modulus error of 0.04 mag (RLGH24). The anchors (/calibrators)222We will call anchors or calibrators to the host galaxies in this first step of our method. used to obtain absolute distances should be galaxies where there is a confluence on the distance estimates obtained by various methods and groups.
Our method does not rely on many host galaxies of SNe Ia, but on a sample of 4 nearby host galaxies for which there is a very good agreement in the distance from the SH0ES and the CCHP collaborations. We use the distances to M101, NGC 7250, NGC 5643, M66 to anchor the twins. The uncertainty in the anchor is a part of the systematic error in the obtained H0, and it varies depending on the anchor from 0 to 0.04 mag. The advantage to select this group of well measured galaxies is to avoid the whole range of discrepancies introduced by the present samples in rung 2 when applying different methods (Cepheids, TRGB, JAGB stars) by the two collaborations. The choice of these galaxies has been driven, as well, by the availability in number of SNe Ia twins in the Hubble flow of the kind found in these low-z galaxies.
The three–rung method (see Figure 6 un Riess et al. 2024) presents a sample of calibrators in rung 2 which looks far from being ideal: the calibrators point to individual H0 values from 60 to 86 km s-1 Mpc-1. The calibrators in rung 2 are taken together with the sample of SNe Ia in the Hubble flow to define a fiducial absolute magnitude MB for SNe Ia and the slope in the Hubble diagram. It seems clear that selecting particular subsamples of these galaxies for calibrating the SNe Ia would produce biases in H0. This fact has been identified as one of the causes of the divergent results for H0 (Riess et al. 2024). In our method, the representation of the types of SNe Ia in the Hubble flow with a given intrinsic luminosity of SN Ia matches the luminosity of its twin SN Ia at low z. The uncertainty will come from the distance of the low z galaxies of each SN Ia type. We will avoid the selection effects arising from modeling a disperse SNe Ia family.
A crucial task now is to verify if the three-rung process, which culminates with distances given to the SNe Ia in the Hubble flow, is in accordance with the distances that we found by our more direct method. Our procedure is then a test of what is the most important delivery of the three rung method: the distances of SNe Ia in the Hubble flow and the value of H0.
The paper is organized as follows: In section 2, we describe the sample of nearby SNe Ia which will have twins in the Hubble flow. In section 3, we present the reasons to use those SNe Ia and their host galaxies as anchors. In section 4, we present our sample of SNe Ia in the Hubble flow. In section 5, we describe our method. In section 6, we present the results on the distances of the SNe Ia. In section 7, we discuss the distance moduli of SNe Ia and the H0 obtained. We compare our results to the distances obtained by the Pantheon+ sample and by the CCHP. In section 8, we present a summary and give our conclusions.
2 The use of twin SNe Ia
In our previous paper (RLGH24), we tested the method in two SNe Ia from the same galaxy. Instead of , we were interested in , the relative distance between the 2 twin SNe Ia, SN 2013aa and SN 2017cbv in NGC 5643. The results point to = 0.00.0, as it should have been expected from the two SNe being in the same galaxy with similar conditions of negligible reddening in the host galaxy and same reddening by dust in our Galaxy, and a 0.004 0.005 in an early spectrum and =–0.023 in a nebular one. Those correspond to a precision of 23 kpc from the early spectrum and around 100 kpc from the late one. The joint result from the two phases gives a distance difference compatible with 0 with -0.005 0.004 mag.
The preceding approach demonstrated the possibility to combine early and late phase information to obtain relative distances or differences in distance moduli between the hosts of SN Ia twins.
To estimate the magnitude accuracy that twins have in their spectral comparison, we calculated the difference in filters B, V, R, I filters of the twin spectra at the same phase once they are shifted to the same distance (this distance is calculated given the information provided by our emcee computation). The error estimate is of 0.04 in all filters and all phases. This confirms the power of this method to disentangle errors in SNe Ia host galaxies distances provided by TRGB, Cepheids, JAGB.
Twin spectra of SNe Ia near maximum should have a similar shape of the pseudo–continuum and similar pseudo–equivalent widths (pWs) of the different lines. All that is required because otherwise it is not possible to have a difference of 0.04 in magnitude in all filters.
Other approaches using the spectral diversity of SNe Ia twins in the context of improving the determination of the nature of dark energy were presented by Fakhouri et al. (2015) and Boone et al. (2016). In our case, we want to use the ”twins” to determine the Hubble constant.
We follow the subclassification of SNe Ia according to the Branch classification scheme (Branch et al. 2006), which has been set in more quantitative terms by the Carnegie Supernova Project. The Branch classification subdivides SNe Ia in four types: Core Normal (CN), Broad Line (BL), Shallow Silicon (SS) and Cool (CL). The main factor to be taken into account is the pseudoequivalent widths (pEW) of the lines in the spectrum. Burrow et al. (2020) show how these classes are well separated in the p(EW) diagrams. The physical explanation of these subclasses is linked to the velocities and chemical composition of the SN Ia core which is revealed as the supernova photosphere recedes showing the inner ejecta. Those differences can be seen as four different subtypes with typical pseudo–equivalent widths (pWs) at various phases.
The spectra of the twins have very similar pseudo-equivalent widths (pWs) in the list of classifying lines of the subtypes (see Morrell et all. 2024). Those lines are pW1 (Ca H & K), pW2 (Si II 4130 Å ), pW3 (Mg 4481; blended with Fe II), pW4 (Fe II at 4600 Å blended SII), pW5 (S II ”W” 5400 Å ), pW6 (Si II 5972 Å ), pW7 (Si II 6355 Å ), pW8 (Ca II IR Å ). In particular, the line associated with Si II 5972 Å should be very similar in twin SNe Ia, with a discrepancy of less than 5. This line correlates with stretch (, sBV), and amongst SNe Ia of similar stretch, has to be almost the same for twins. In the pW6 over pW7 diagram, the twins should fall in very close positions. If the equivalent widths are similar, the ratios of the two Si II at 5972 Å and 6355 Å , , should coincide.
We looked, in our method paper (RLGH24), at the abovementioned signs for the classification as twins of nearby SNe Ia.
In that previous paper (RLGH24), we used a single subtype of SNe Ia: the Core Normal (CN) subtype. In the present work we have extended it to the Broad Line (BL) subtype. We have found a good anchor of BL SNe Ia: SN 1989B in M66.
Previous research done within the core normal subtype started with a comparison between the twins SN 2013aa/SN 2017cbv, SN 2013aa and SN 2017cbv appeared in the outskirts of the same host galaxy, NGC 5643, so the reddening E(B-V) should be negligible. The reddening in our Galaxy in the direction of NGC 5643 is = 0.15 mag (Schlafly & Finkbeiner, 2011). SN 2013aa and SN 2017cbv have similar decline rates and B-peak magnitudes. The studies of these two SNe Ia also reveal similar characteristics in other aspects. Table 1 specifies the decline rates of those SNe Ia not only by but also by s (Burns et al. 2020). Fitting the B-V color curves with cubic splines, Burns et al (2020) find identical color-stretch s = 1.11 for SN 2013aa and SN 2017cbv. This value is smaller (slower decline) than in typical SNe Ia. They also found similar values in the Branch classification, concerning the pseudo equivalent widths (pEW) of the lines of different elements. This proves that their interior was very alike in chemical composition. The use as anchor of the host galaxy of both SN 2013aa and SN 2017cbv NGC 5643 is justified given the converging values of distance moduli to this galaxy with the use of Cepheids by SH0ES and with JAGB stars by the CCHP.
In the previous paper, we found that SN 2013dy is practically equal at late phases to SN 2017cbv (there are no spectra of SN 2013aa at that phase to allow a comparison). A comparison shows that the light curve at late phases is slightly more luminous than in SN 2013aa, but coicncides with SN 2017cbv. At very early phases it is similar to both SN 2017cbv and SN 2013aa though slightly redder. After maximum brightness, it is also identical to SN 2013aa. SN 2013dy has a significant reddening. The reddening in the Galaxy is E(B-V)MW 0.14 mag (Schlafly & Finkbeiner 2011), though there are indications that the total reddening might be 0.15 mag. There is the possibility, then, to use SN 2013dy as an anchor. The fact that this supernova was discovered a few hours after explosion and had a very intense photometric and spectroscopic follow up (Pan et al. 2015) enables to have many phases for comparison with their SNe Ia twins in the Hubble flow. There are, as well, converging values of distance moduli to the host galaxy of this SNIa, NGC 7250, with the use of Cepheids by SH0ES and with JAGB stars by the CCHP.
SN 2011fe, of the same Branch subtype as SN 2013aa, SN 2017cbv, and SN 2013dy, is also identified as a very good reference as both SH0ES and CCHP find similar distances to the host galaxy M101. SN 2011fe was discovered in its very early phase and was classified as a normal SN Ia (Nugent et al. 2011). At maximum light, the color was (Bmax - Vmax) = –0.07 0.02 mag. The value of (Bmax - Vmax) being –0.12 mag in normal SNe Ia, suggests that there is some host galaxy extinction.
Concerning the luminosity decline parameter of SN 2011fe, Burns et al (2023, private communication) measured a = 1.070.006 mag.
SN 2011fe is a faster and more underluminous SN Ia than SN 2017cbv and SN 2013aa. The difference between them is seen at both early and late phases. In fact, SN 2013aa and SN 2017cbv synthesized about 0.23 M⊙ more 56Ni than SN 2011fe. Studies by Jacobson-Galán et al. (2018) suggest that SN 2013aa and SN 2017cbv synthezised 0.732 0.151 M⊙ of 56Ni.
| SN 2013aa | ||
| RA, DECa | 14:32:33.881 | -44:13:27.80 |
| Discovery datea | 2013-02-13 | |
| Phase ( referred to maximum light)b | -7 days | |
| Redshiftc | 0.004 | |
| E(B-V)MWd | 0.150.06 mag | |
| b | 11.0940.003 mag | |
| b | 0.950.01 | |
| Stretch factor b | 1.110.02 | |
| Phases of the spectra used | -2, 34, 42.07 days | |
| SN 2017cbv | ||
| RA, DECe | 14:32:34.420 | -44:08:02.74 |
| Discovery datee | 2017-03-10 | |
| Phase (referred to maximum light)b | -18 days | |
| Redshiftc | 0.004 | |
| E(B-V)MWd | 0.150.06 mag | |
| b | 11.110.03 mag | |
| b | 0.960.02 | |
| Stretch factor b | 1.110.03 | |
| Phases of the spectra used | 13.14 days | |
| SN 2013dy | ||
| RA, DECf | 22:18:17.599 | +40:34:09.59 |
| Discovery datef | 2013-07-10 | |
| Phase (referred to maximum light)g | -18 days | |
| Redshifth | 0.0039 | |
| E(B-V)MWd | 0.15 mag | |
| i | 13.2290.010 mag | |
| j | 0.920.006 mag | |
| Stretch factor j | 1.0910.03 | |
| Phases of the spectra used | 3.7, 5 days | |
| SN 2011fe | ||
| RA, DECa | 14:03:05.810 | +54:16:25.39 |
| Discovery datek | 2011-08-24 | |
| Phase (referred to maximum light)l | -20 days | |
| Redshiftm | 0.0012 | |
| E(B-V)MWd | 0.008 mag | |
| n | 9.9830.015 mag | |
| 1.070.06 | ||
| Stretch factor | 0.97 0.002 | |
| Phases of the spectra used | 4.65, 13 days | |
| SN 1989B | ||
| RA, DECo | 11:20:13.900 | +13:00:19.00 |
| Discovery dateo | 1989-01-30 | |
| Phase (referred to maximum light)p | -7 days | |
| Redshiftq | 0.00242 | |
| E(B-V)MWp | 0.32 mag | |
| q | 1.0530.11 mag | |
| Stretch factor q | 0.9540.034 mag | |
| Phases of the spectra used | -1, 3, 7, 9, 11 days |
| aWaagen (2013). | bBurns et al. (2020). |
| cParrent et al. (2013). | dSchlafly & Finkbeiner (2011). |
| eTartaglia et al. (2017). | fCasper et al. (2013). |
| gZheng et al. (2013). | hSchneider et al. (1992). |
| iPan et al. (2015). | jUddin et al (2024) |
| kRichmond & Smith (2012) | lCenko et al. (2011) |
| mBurns (private communication, 2023). | nYang et al. 2020. |
| pEvans (1989). | qMorrell et al. (2024). |
| rWells et al. (1994). |
These four SNe Ia are the anchors for the CN SNe Ia in the Hubble flow. For BL SNe Ia in the Hubble flow, we can use as anchor SN 1989B. This supernova was studied in detail by Wells et al. (1994). We have been able to compare with our method significantly reddened SNe Ia in the Hubble flow with SN 1989B. One of the advantages of SN 1989B is that the distance determination coincides using the Cepheids (Saha et al. 1999) and the TRGB by Freedman et al (2019).
We present these SNe Ia in Table 1.
3 On the convergence of the distance to our anchors
3.1 M101
Recently, a convergence has been achieved on the distance to M101 by the SH0ES group and the CCHP. This gives an extraordinary opportunity to be safe on the class of SN 2011fe–like supernovae. The distance to M101 is 29.188 0.055 mag for the SH0ES group using Cepheids and 29.18 0.04 mag also from Cepheids by the CCHP in 2024 (Freedman et al. 2024) and 29.151 0.04 mag using TRGBs. In the most recent CCHP publication of the distance to this galaxy (Freedman et al. 2025), it is only slightly different.
There are other approaches giving slightly smaller distances to M101, like the Mira variables (Huang et al. 2024) and the Blue Supergiants (Bresolin et al 2025). All are within 1 of the values from the SH0ES collaboration and the latest CCHP value quoted above. The JAGB stars distance to M101 from the CCHP (Freedman et al. 2025) is 29.208 0.045 mag.
3.2 NGC 5643
In the most recent papers by the SH0ES collaboration (years 2024–2025), the distances to NGC 5643 are given not only using the Cepheids method, but also the other methods used by the CCHP.
The values from SH0ES in the Table 2 from Li et al. (2024) include TRGB distances obtained with the JWST and they are compared with their Cepheids distance determination to the same galaxies. Their TRGB values are not far from the Cepheids ones (only differing by 0.02 mag at most). The Cepheid distance modulus of reference for NGC 5643 is 30.55 0.063 mag, which is the same as in Riess et al. (2022). Their TRGB values slightly change if one uses different values for their smoothing s 0.10 or s 0.05. Fortunately, for NGC 5643 the difference is small ( 0.01): the distance is 30.57 0.06 mag, when using a smoothing of s=0.10, or 30.58 0.06 mag when using a smoothing of s=0.05. The latest value from the TRGB method for NGC 5643 in 2025 obtained by the CCHP collaboration is 30.643 0.066 mag (Freedman et al. 2025). This value has changed often during the last year. The distance has gone in the last months from 30.61 0.07 mag (Freedman et al. 2024) to 30.643 0.066 mag (Freedman et.al. 2025). There is a considerable difference in the TRGB values provided by SH0ES and by the CCHP.
Fortunately, there is an agreement on the distance to NGC 5643 between Cepheids used by SH0ES (Riess et al. 2022) and JAGB stars used by the CCHP (Freedman et al. 2025). As mentioned above the Cepheid distance value by SH0ES is 30.55 0.06 mag, and their TRGB value 30.57/30.58 0.06 mag. Whereas the TRGB value by the CCHP has been changing often, their JAGB value seems stable at 30.582 0.038 mag. (The JAGB value by SH0ES is 30.54 0.04, well in line with the previous values). We compare then our results using as anchor for this galaxy the Cepheids by SH0ES with the value anchored in the JAGB stars by the CCHP in section 7.
3.3 NGC 7250
In NGC 7250 there is good agreement between the distance obtained by SH0ES and the distance obtained by the CCHP with the JAGB stars.
The latest SH0ES measurement gives a distance of 31.49 0.05 mag using Cepheids measured with the JWST (Appendix 1 in Riess et al. 2025). The distance to NGC 7250 with Cepheids before the JWST had a 0.125 mag error. There is a very good agreement between the SH0ES collaboration and the CCHP collaboration using JAGB stars. The value by the SH0ES project is 31.59 0.04 mag by Li et al. (2025) and 31.59 0.02 (stat) 0.04 (sys) mag by the the CCHP project (Lee et al. 2025). This makes of NGC 7250 a good anchor, as central values coincide. On the contrary, the value using TRGB by the CCHP is 31.629 0.034 and it is different from the TRGB by SH0ES.
3.4 M66
This is a crucial SN Ia host whose distance has been been measured with TRGB by Freedman et al. (2019) giving a distance modulus of 30.22 0.04 mag, which coincides with the distance modulus obtained by Saha et al. (1999) of 30.22 0.12 mag. It looks like a good anchor for the BL SNe Ia in the Hubble flow. Both measurements are based on images taken with the HST. It is ideal for getting distances to SN 2011cn in IC 4758, SN 2008go in 2MASX J22104396-2047256, SN 1999ek in UGC 3329.
For the comparison of SN 1989B and the SNe Ia which are BL but also high velocity (HV), we need to apply a shift to the reference SN 1989B in accordance with the HV they have.
4 Sample of SNe Ia in the Hubble flow
The selection of twins has to follow the requirements of being within the same subtype of SNe Ia (within a given Branch et al. 2005 class), have a similar stretch within errors, and having various epochs for the spectra. A number of SNe Ia from the Carnegie Supernova Project (CSP) fullfil those criteria.
Table 2 shows the references for the spectra and photometry used in this paper. Some spectra come directly from the Carnegie Supernova Project archive, and some others taken from the WISeREP. All are carefully calibrated in flux in agreement with the published photometry. The error in the flux calibration is taken into account.
The comparison of the light curves of the Hubble flow SNe Ia with those nearby, is a first selection criterion in the use of the ”SNe Ia twins for life” method. We have taken similar and the stretch factor (see Tables 3 to 14). Burns et al. (2014) describe a relation between the SNooPy stretch and , which we find as an approximation:
where is the decline rate parameter established by Phillips (1993), which is at the core of the discovery of the correlation of the brightness of a SN Ia and the rate of decline of the B light curve in the 15 days past maximum. The color stretch is the rate of decline of the B-V color in 30 days. One disadvantage of is that one needs to have a photometric B and V coverage from B maximum until approximately 40 days past maximum. The CSP and Burns et al. (2020) quote, instead of , the stretch , where the “D” in the superindex refers to the direct comparison in B, without having to take into account the BVRI light curves,
In this research, we have been very careful in selecting for the comparison of SNe Ia in the Hubble flow with their twins, SNe Ia with very similar and stretch values.
| Supernova | Galaxy | Spectral range (Å) | Phase | Date | Reference | |
| SN 2012bo | NGC 4726 | 0.025431 | 3665-8847 | +14 | 2012-04-18 | WISeREP |
| 3634-9602 | +26 | 2012-04-30 | ||||
| SN 2008bq | ESO 308-G25 | 0.03400 | 3675-8931 | +5.57 | 2008-04-12 | CSP-DR1 |
| +33.44 | 2008-05-10 | |||||
| SN 2008bz | A123857+1107 | 0.0603 | 3626-9433 | +4.65 | 2008-04-26 | |
| 3800-9235 | +13 | 2008-05-05 | ||||
| LSQ12fxd | ESO 487-G004 | 0.0312 | 3352-7475 | -2 | 2012-11-13 | CSP-DR1 |
| 3579-9615 | +3.7 | 2012-11-19 | ||||
| SN 2008bf | NGC 4055/4057 | 0.0235 | 3715-9974 | -2.70 | 2008-03-26 | |
| 3710-9018 | +42.07 | 2008-05-11 | ||||
| SN 2007A | NGC 105 | 0.01764 | 3473-7409 | -2.86 | 2007-01-10 | WISeREP |
| 3378-1027 | +13.14 | 2007-02-24 | ||||
| ASASSN-15db | NGC 5996 | 0.01099 | 3604-9164 | +3 | 2015-02-25 | |
| +4 | 2015-02-26 | |||||
| LSQ14gov | GALEXMSC J040601.67-160139.7 | 0.0896 | 3650-9250 | -6 | 2013-12-30 | |
| 3430-9100 | +1 | 2014-01-06 | ||||
| SN 2007ca | MCG -02-34-61 | 0.01407 | 3455-9494 | +1.4 | 2007-05-08 | CSP-DR1 |
| 3297-9416 | +4.2 | 2007-05-11 | ||||
| SN 2001cn | IC 4758 | 0.0156 | 3827-7218 | +2.2 | 2001-06-14 | |
| 3715-7115 | +10 | 2001-06-22 | ||||
| SN 2008go | 2MASX J22104396-2047256 | 0.0623 | 3800-9234 | +0.6 | 2008-10-27 | |
| 3636-9450 | +7.5 | 2008-11-04 | ||||
| SN 1999ek | UGC 3329 | 0.0177 | 3720-7540 | +3.3 | 1999-11-03 | |
| 3720-7540 | +9.32 | 1999-11-09 | ||||
| Photometry | ||||||
| SN 2012bo | Morrell | |||||
| SN 2008bq | CSP-DR1 | |||||
| SN 2008bz | CSP-DR1 | |||||
| LSQ12fxd | Morrell | |||||
| SN 2008bf | CSP-DR1 | |||||
| SN 2007A | CSP-DR1 | |||||
| ASASSN-15db | Morrell | |||||
| LSQ14gov | Morrell | |||||
| SN 2007ca | CSP-DR1 | |||||
| SN 2001cn | Krisciunas | |||||
| SN 2008go | CSP-DR1 | |||||
| SN 1999ek | Krisciunas |
| SN 2012bo in NGC 4726 | ||
| RA, DECa | 12:50:45.23 | -14:16:08.5 |
| Discovery datea | 2012-03-27 | |
| Phase (referred to maximum light)b | -8 days | |
| Redshiftb | 0.026548 | |
| E(B-V)MWc | 0.046 mag | |
| Stretch factor b | 1.1600.0411 | |
| b | 0.9310.061 | |
| Phases of the spectra used | 14, 26 days |
| aItagaki et al. (2012). | bMorrell et al. (2023). |
| cSchlafly & Finkbeiner (2011). |
| SN 2008bq in ESO 308-G25 | ||
| RA, DECa | 06:41:02.51 | -38:02:19.0 |
| Discovery datea | 2008-04-02 | |
| Phase (referred to maximum light)b | -4 days | |
| Redshiftb | 0.03446 | |
| E(B-V)MWc | 0.080 mag | |
| Stretch factor d | 1.0560.014 | |
| d | 0.8950.062 | |
| Phases of the spectra used | 5.57, 33.44 days |
| aLuckas et al. (2008). | bFolatelli et al. (2013). |
| cSchlafly & Finkbeiner (2011). | dKrisciunas et al. (2017). |
| SN 2008bz in A123957+1107 | ||
| RA, DECa | 12:38:57.74 | +11:07:46.2 |
| Discovery datea | 2008-04-22 | |
| Phase (referred to maximum light)b | 0 days | |
| Redshiftb | 0.0614 | |
| E(B-V)MWc | 0.023 mag | |
| Stretch factor d | 0.9480.046 | |
| b | 1.0930.079 | |
| Phases of the spectra used | 4.65, 13 days |
| aYuan et al. (2008). | bKrisciunas et al. (2020). |
| cSchlafly & Finkbeiner (2011)t al. (2023). | dMorrell et al. (2023). |
| LSQ12fxd in ESO487-G004 | ||
| RA, DECa | 05:22:17.02 | -25:35:47.1 |
| Discovery datea | 2012-10-31 | |
| Phase (referred to maximum light)b | -15 days | |
| Redshiftb | 0.031 | |
| E(B-V)MWc | 0.022 mag | |
| Stretch factor d | 1.1610.040 | |
| b | 0.9400.060 | |
| Phases of the spectra used | -2, 3.7 days |
| aBufano et al. (2012). | bMorrell et al. (2023). |
| cSchlafly & Finkbeiner (2011). | dMorrel et al. (2023). |
| SN 2008bf in NGC 4055/4057 | ||
| RA, DECa | 12:04:02.90 | +20:14:42.6 |
| Discovery datea | 2008-03-18 | |
| Phase (referred to maximum light)b | -11 days | |
| Redshiftb | 0.02510 | |
| E(B-V)MWc | 0.036 mag | |
| b | 15.7790.022 mag | |
| Stretch factor d | 1.0580.012 | |
| d | 0.970.060 | |
| Phases of the spectra used | -2.70, 42.07 days |
| aPanski et al.(2008). | bFolatelli et al. (2013). |
| cSchlafly & Finkbeiner (2011). | dKrisciunas et al. (2017). |
| SN 2007A in NGC 105 | ||
| RA, DECa | 00:25:16.66 | +12:53:12.5 |
| Discovery datea | 2007-01-02 | |
| Phase (referred to maximum light)b | -10 days | |
| Redshiftb | 0.01648 | |
| E(B-V)MWc | 0.074 mag | |
| Stretch factor d | 1.0120.061 | |
| d | 0.9710.074 | |
| Phases of the spectra used | -2.86, 13.14 days |
| aPicket et al. (2007). | bFolatelli et al. (2013). |
| cSchlafly & Finkbeiner (2011). | dKrisciunas et al. (2017). |
| ASASSN-15db in NGC 5996 | ||
| RA, DECa | 15:46:58.9 | 17:53:03 |
| Discovery datea | 2015-02-15 | |
| Phase (referred to maximum light)b | -7 days | |
| Redshiftb | 0.0114 | |
| E(B-V)MWc | 0.03 mag | |
| Stretch factor b | 0.9530.040 | |
| b | 1.0890.061 | |
| Phases of the spectra used | 3, 4 days |
| a Holoien et al. (2015). | bMorrell et al. (2023). |
| cSchlafly & Finkbeiner (2011). |
| LSQ14gov in GALEXMSC J040601.67-160149.7 | ||
| RA, DECa | 04:06:01.53 | -16:01:38.86 |
| Discovery datea | 2014-12-21 | |
| Phase (referred to maximum light)b | -15 days | |
| Redshiftb | 0.0896 | |
| E(B-V)MWc | 0.0377 mag | |
| Stretch factor b | 1.1410.041 | |
| b | 0.9790.061 | |
| Phases of the spectra used | -6, 1 days |
| aDucrof et al. (2014). | b Morrell et al. (2023). |
| cSchlafly & Finkbeiner (2011). |
| SN 2007ca in MGC-02-34-61 | ||
| RA, DECa | 13:31:05.81 | -15:06:06.6 |
| Discovery datea | 2007-04-25 | |
| Phase (referred to maximum light)b | -12 days | |
| Redshiftb | 0.01508 | |
| E(B-V)MWb | 0.067 | |
| Stretch factor b | 1.0770.012 | |
| b | 0.9040.032 | |
| Phases of the spectra used | 1.4, 4.2 days |
| aItagaki et al. (2007). | bUddin et al. (2020). |
| SN 2001cn in IC 4758 | ||
| RA, DECa | 18:46:17.84 | -65:45:41.8 |
| Discovery datea | 2001-06-12 | |
| Phase (referred to maximum light)b | -4 days | |
| Redshiftb | 0.0156 | |
| E(B-V)MWc | 0.052 | |
| Stretch factor b | 0.9270.020 | |
| b | 1.0800.044 | |
| Phases of the spectra used | 2.2, 10 days |
| aChassagne (2001) | bMorrell et al. (2024) |
| cSchlafly & Finkbeiner (2011). |
| SN 2008go in 2MASX J22104396-2047256 | ||
| RA, DECa | 22:10:44.03 | -20:47:17.2 |
| Discovery datea | 2008-10-22 | |
| Phase (referred to maximum light)b | -5 days | |
| Redshiftc | 0.06 | |
| E(B-V)MWb | 0.033 | |
| Stretch factor b | 0.9390.044 | |
| b | 1.0930.089 | |
| Phases of the spectra used | 0.6, 7.5 days |
| aGriffith et al. (2008). | bStahl et al. (2019). |
| cSchlafly & Finkbeiner (2011). |
| SN 1999ek in UGC 3329 | ||
| RA, DECa | 05:36:31.60 | +16:38:17.8 |
| Discovery datea | 1999-10-20 | |
| Phase (referred to maximum light)b | -11 days | |
| Redshiftb | 0.0177 | |
| E(B-V)MWb | 0.502 | |
| Stretch factor b | 0.9210.008 | |
| b | 1.0970.019 | |
| Phases of the spectra used | 3.3, 9.32 days |
| aJohnson & Li (1999. | bMorrell et al. (2024). |
4.1 The SN 2011fe–like in the Hubble flow
We have thus M101 and its supernova SN 2011fe at a distance of d 6.880 0.174 Mpc ( 29.188 0.055). In the Hubble flow we have found an excellent twin in SN 2008bz. It has a of 1.093 0.079. The of SN 2011fe is 1.07 0.006.333The two SN 2011fe-like SNe Ia have the following stretches: SN 2008bz, = 0.9480.9480.046; ASASSN-15db, = 0.9530.04. Both compare well with SN 2011fe, with = 0.970.033. SN 2008bz’s heliocentric redshift is 0.06, thus the full expansion of the modulus in q0 is needed. The result can be seen in Figure 1. It is remarkable that with this method we can go from 6.8 Mpc to 270 Mpc in a single step. We will come back to the result in distance in section 7.
Another SN Ia in the Hubble flow of the same type as SN 2011fe is ASASSN-15db in NGC 5996. It has a nearly identical light–curve decline values as SN 2011fe, with 1.089 0.061 for ASSASN-15db and 1.07 0.06 for SN 2011fe. Despite the fact that we do not have the full light curve of this SN Ia (which belongs to the DR of CSP II soon to come), we can calibrate the spectra with the photometry for the two phases.
4.2 SNe Ia in the Hubble flow like SN 2013aa/2017cbv
SNe Ia that can be considered twins of SN 2013aa are SN 2007A, SN 2008bf and LSQ14gov.
SN 2007A has 0.9710.074 which compares well with SN 2013aa and SN 2017cbv, since 0.950.01 for SN 2013aa.444The stretches of the three SNe Ia of this group are: for SN 2007A, = 1.0250.061; for SN 2008bf, = 1.0580.012; for LSQ14gov, = 1.1410.041. All of them compare well with the stretch of the anchor SN 2013aa: = 1.110.02.
This supernova has an internal reddening of 0.2 magnitudes. Such a high extinction by dust has been well captured by the SNooPY light curve fitter. It was not well captured by SALT2 and, therefore, the value of the distance derived by Pantheon+ to this supernova is too large.
SN 2008bf compares very well with SN 2013aa. The first one has a 0.97 0.06 and the second one 0.950.01. For the case of doubt about their similarity, a direct comparison of the light curves can be seen for this one and for all SNe Ia from CSP I (see Appendix). SN 2008bf as a perfect twin of SN 2013aa.
For the last comparison, we have LSQ14gov which has 0.9790.061 to compare with 0.95 0.01 of SN 2013aa.
4.3 SNe Ia in the Hubble flow like SN 2013dy
The advantage of these SNe Ia is that there is a pretty good agreement between the distances given by the two groups for this anchor.
Four SNe Ia from the SN 2013dy class are SN 2012bo and SN 2008bq, SN 2007ca and LSQ12fxd.
SN 2012bo has a 0.9310.061 which compares well with the 0.92 0.006 from SN 2013dy.555The stretches of the four SNe Ia of this group are: for SN 2012bo, = 1.1600.0411; for SN 2008bq, = 1.0560.014; for SN 2007ca, =1.0770.012; for LSQ12fxd, = 1.1610.040. All of them compare well with the stretch of the anchor SN 2013dy: = 1.0910.03.
We have paired SN 2008bq with SN 2013dy. It has a similar of 0.90 0.062 versus 0.920.006 for SN 2013dy.
SN 2007ca has 0.9040.032 while is 0.92 0.006 for SN 2013dy. They are very good twins.
LSQ12fxd has 0.9400.06, versus 0.920.006 of SN 2013dy.
4.4 SNe Ia in the Hubble flow like SN 1989B
SN 1989B is a broadline SNe Ia that appeared in M66, a nearby galaxy for which there is a coincident distance by TRGBs (Freedman et al. 2019) and Cepheids (Saha et al. 1999). There is no published distance by SH0ES.
Three SNe Ia in the CSP I are broad line and have a rate of decline in the light curve similar to SN 1989B. Usually broadline supernovae have also high velocity ejecta. This produces a blueshift in the lines of the supernova. SN 1989B is not a high velocity supernova. Two of the broadline SNe Ia at high–z have high velocity features: SN 2001cn and SN 2008go. In contrast, SN 1999ek is a broad line that does not exhibit high velocity in the its spectral features.
We have paired SN 2001cn, SN 2008go and SN 1999ek with SN 1989B. There will be a blueshift applied to SN 1989B to compare with SN 2001cn and SN 2008go, whereas in SN 1999ek it is not necessary.
SN 2001cn has = 1.0800.044, which fits well with = 1.0530.11 of SN 1989B.
SN 2008go has = 1.0930.089, also concordant with the = 1.0530.11 of SN 1989B.
SN 1999ek has = 1.0970.019, again compatible with the = 1.0530.11 of the anchor666Concerning the stretch, SN 1989B has = 0.9540.034. The corresponding values for its distant twins are: for SN 2001cn, = 0.9270.020; for SN 2008go, = 0.9390.044; for SN 1999ek, = 0.9210.008. So, the stretches of the three distant SNe Ia compare well with that of the anchor SN !989B.
5 Methodology
We have used eight SNe Ia from Carnegie I included in the DR1. Their characteristics are described in Folatelli et al. (2012) and four SNe Ia in Carnegie II (Morrell et al. 2024). We have been able to calibrate the spectra, which are uploaded to the WISeREP.
All the spectra have been corrected from Galactic reddening. This is done with the purpose of allowing to compare Galactic dust-free spectra. It also helps to learn about dust extinction in the host. The spectra are compared in the frame of the Hubble flow supernova. For that purpose nearby spectra are shifted in wavelength and their flux is diluted in accordance with the (1 + ) law. The shift brings the nearby SN Ia to the redshift of the Hubble flow supernova taking into account the heliocentric redshift of the nearby SNIa (shifting by the difference, though the redshift of nearby SNe Ia is always low).
The method primarly determines the distance between the nearby SN Ia and that in the Hubble flow. It will also give the relative reddening. The value will indicate whether is larger or smaller for one of the two SNe with respect to its twin, taken as reference. Concerning the distance modulus, in reality we are calculating distances relative to the SNe Ia of reference. So, we have a distance scale factor between the reference and the SN in the Hubble flow. For the sake of producing a final result, we give the absolute distance modulus .
To find the best values and the uncertainties of all variables, we explore the parameter space with Markov Chain Monte Carlo (MCMC) techniques after converting the into a log likelihood function (the probability of a dataset given the model parameters) that we aim to optimize:
where obs corresponds to a given supernova and ref to its reference twin, which represents a whole class. and correspond to the fluxes of the two SNe at different wavelengths, and includes the quadratic addition of the uncertainty on the fluxes of both SNe. The MCMC allows to get the maximum likelihood values of each parameter from the posterior distribution of samples, together with their uncertainties, including the propagation of all the parameter uncertainties on each parameter.
In fact, as we are using different phases, we have a total likelihood function. For two phases, corresponds to 1 and 2 below. Thus the total likelihood function is written as
i.e. just the addition of all individual SN phases.
We use a Python package, EMCEE (Foreman–Mackey et al. 2013)777https://emcee.readthedocs.io/ to explore the likelihood of each variable. EMCEE utilizes an affine invariant MCMC ensemble sampler proposed by Goodman et al. (2010). This sampler tunes only two parameters to get the desired output: number of walkers and number of steps. The run starts by assigning initial maximum likelihood values of the variables to the walkers. The walkers then start wandering and explore the full posterior distribution. After an initial run, we inspect the samplers for their performance. We do this by looking at the time series of variables in the chain and computing the autocorrelation time, 888https://emcee.readthedocs.io /en/stable/tutorials/autocorr/. In our case, the maximum autocorrelation time among the different variables is about 50. When the chains are sufficiently burnt-in, they forget their initial start point), we can safely throw away some steps that are a few times higher than the burnt-in steps. In our case, we run EMCEE with 32 walkers and 10,000 steps and throw away the first 250 samples, equivalent to 4 times the maximum autocorrelation time. Thus, our burn–in value in this computation is equal to 4 times the maximum autocorrelation time. A criterion of good sampling is the acceptance fraction, . This is the fraction of steps that are accepted after the sampling is done. The suggested value of is between 0.2 - 0.5 (Gelman et al. 1996). In each run, we typically obtain 0.25.
One can visualize the output of two-dimensional and one-dimensional posterior probability distributions in a corner plot corresponding to 1, 2, 3. This corner plot shows one and two dimensional projections of the posterior probability distributions of our parameters.
6 Distances of the SNe Ia in the Hubble flow
We present here the results of the MCMC for all the SNe Ia in the Hubble flow that have been included in this paper. We give the results for each phase and the combined phase (see Figures 1 to 12).
7 Results and discussion
7.1 The test of the three rung method
The present work can serve to test the three rung method for the determination of H0. In this respect, the slope obtained by the SH0ES-Pantheon+ collaboration and by our method using only one step is the same. That could in principle not have been the case, as the 3 rung method goes through the calibration of a fiducial absolute magnitude of SNe Ia which is not required by our method.
We include in Table 15 the distance moduli using the ”SNe Ia twins for life” method as well as those provided by Pantheon+ with SH0ES and by the Carnegie Supernova Project. Those last two methods are calibrated with the Cepheid distances. The SNe Ia moduli in the Hubble flow from the Pantheon+ and SH0ES are quite close to those from the CCHP.
The Pantheon+ data come with their errors in distance moduli in their release and the CCHP distance moduli and errors are generated as in Uddin et al. (2024) code given the entries in the light curve parameterization. We see similar distances moduli in Table 15 by our method when compared with those from Pantheon+ and with Cepheids by the CCHP. There are some differences (in 3 SNe Ia) due to the underestimate of reddening (see Figure 13), as we will discuss in the next subsection.
| This work | SH0ES | CCHP | ||||
| SN | ||||||
| 2012 bo | 35.1630.0928 | 75.2334.287 | 35.446 0.217 | 66.0406.599 | ||
| 2008bq | 35.8260.0898 | 72.4523.661 | 35.7280.178 | 75.7965.964 | 35.8790.205 | 70.7056.893 |
| 2008bz | 37.2150.0777 | 69.3902.728 | 37.1680.288 | 71.0723.945 | 37.1550.280 | 71.3349.198 |
| LSQ12fxd | 35.5510.0908 | 74.6663.928 | — | — | - 35.6150.207 | 72.4976.911 |
| 2008bf | 35.0710.0879 | 74.1154.211 | 35.0520.231 | 74.7167.494 | 35.0110.222 | 76.1927.789 |
| 2007A | 34.2650.0995 | 70.2675.435 | 34.5340.342 | 62.0439.041 | 34.3170.312 | 68.6049.857 |
| ASASSN-15db | 33.5310.0848 | 68.0916.513 | — | — | 33.5320.330 | 67.75710.297 |
| LSQ14gov | 37.9150.085 | 74.586 3.037 | — | — | 37.9400.185 | 73.7326.336 |
| SN 2007ca | 33.942 0.087 | 74.371 5.765 | 34.178 0.396 | 66.709 13.378 | 34.344 0.278 | 61.802 7.912 |
| This work | Cepheids | CCHP | ||||
| SN | ||||||
| 2001cn | 33.948 0.0661 | 74.2945.421 | 33.8870.2949 | 76.421 11.119 | — | — |
| 2008go | 37.0860.0.067 | 73.561 2.568 | 36.96 0.19 | 77.955 7.128 | 37.004 0.193 | 76.392 6.790 |
| 1999ek | 34.331 0.069 | 72.977 4.743 | — | — | — | — |
To obtain the error in the ”SNe Ia twins” distance modulus we add the various uncertainty terms. The most important one is the systematic error due to different values in the distance of the anchor of the very nearby twin supernovae. As previously said, these anchors converge in value. In the above calculations, errors are of the order of 0–0.04 mag. For broadline SNe Ia, the distance to M66 is the same with the use of TRGB by the CCHP (Freedman et al. 2019) and with Cepheids (Saha et al. 1999). It is at 20.22 0.04 mag. The twins of SN 1989B in the Hubble flow provide a good insight into the value of H0 because of this perfect coincidence between the Cepheids and the TRGB methods by the two collaborations.
We have modeled the errors in the distance moduli for our method adding the various contributionss. We have added in quadrature (see 4) the error in the fit of the distance modulus obtained with the MCMC, , and the covariance between the modulus and the (E(B-V)), cov (, E(B-V)), which is very low for our supernovae. We have as well the error due to a difference of stretch between the Hubble flow SNeIa and their nearby twins that goes into the distance modulus . We model it taking into account the CSP impact of the error in stretch in the distance moduli of the SNe Ia. We must say that, in our method, twins at low z and in the Hubble flow have very close stretches: within 0.02 in sBV. The light curves agree so well between low z SNe Ia and their twins in the Hubble flow (once one makes the cosmological (1+z) time broadening and magnitude dilution) that, in their plot of one over the other, they are often indistinguishable (see Appendix).
The error budget in is:
where is the difference of stretch between the SNe Ia in the Hubble flow and in the low z twins.
For the H0 error calculation, we incorporate the error in peculiar velocity affecting the cosmological z value. This error in the peculiar velocity is not attributed to the distance modulus because it enters into the H0 calculation rather than in the distance. We adopt an error of 400 km s-1 in the calculation of the cosmological redshift (the cosmological redshifts are adopted from NASA/IPAC Extragalactic Database). This error on H0 and those from the terms that enter in (5, see below) are added to the error budget for H0.
We obtain the individual values of H0 for each measured SN Ia.
To calculate the value of H0 we use the expression valid for a flat cosmology, 0:
where H0 is the Hubble constant, D is the distance to the SN Ia, c is the speed of light, z is the cosmological redshift, zHe the heliocentric redshift. The first and second order terms multiplied by z in the parenthesis involve the deceleration parameter, q0, and the cosmic jerk, j0. q0 which taking 0.3 give q0 –0.55 ( 1 - for a flat cosmology). The cosmic jerk can be taken to be j0 1 for the range of redshifts that we are using.
| Supernova | ||||
| SN 2012bo | 0.046 | 0.0480.004 | - | 0.0120.006 |
| SN 2008bq | 0.060 | 0.0670.002 | 0.0130.0255 | 0.0450.009 |
| SN 2008bz | 0.023 | 0.0480.002 | -0.1360.0235 | -0.0060.006 |
| -0.1370.03665 | ||||
| LSQ12fxd | 0.022 | 0.0250.003 | - | -0.0060.006 |
| SN 2008bf | 0.036 | 0.0470.006 | -0.1490.0225 | -0.0460.006 |
| -0.1490.02657 | ||||
| -0.0560.02664 | ||||
| SN 2007A | 0.074 | 0.2830.012 | 0.1170.0255 | 0.2010.010 |
| 0.0630.03365 | ||||
| ASASSN-15db | 0.03 | 0.1080.003 | - | 0.0290.003 |
| LSQ14gov | 0.0377 | 0.04270.001 | - | -0.1040.008 |
| SN 2007ca | 0.067 | 0.3000.001 | 0.2040.0245 | 0.2600.010 |
| 0.1510.02357 | ||||
| 0.2070.02864 | ||||
| SN 2001cn | 0.052 | 0.1800.003 | 0.1180.027 | - |
| SN 2008go | 0.033 | 0.0400.003 | -0.0160.0235 | 0.0420.010 |
| -0.0070.03551 | ||||
| SN 1999ek | 0.502 | 0.6900.006 | - | - |
| 5CSP 50LOWZ/JRK07 51LOSS1 57LOSS2 64CfA3K 65CFA4p2 | ||||
| SN | Distance factor | Anchor |
| 2012bo | 5.183 0.016 | 2013dy |
| 2008bq | 7.033 0.018 | 2013dy |
| 2008bz | 40.315 0.125 | 2011fe |
| LSQ12fxd | 6.196 0.025 | 2013dy |
| 2008bf | 8.019 0.082 | 2013aa |
| 2007A | 5.533 0.010 | 2013aa |
| ASASSN-15db | 7.391 0.049 | 2011fe |
| LSQ14gov | 29.722 0.070 | 2013aa |
| SN2007ca | 2.954 0.003 | 2013dy |
| SN 2001cn | 5.567 0.027 | 1989B |
| SN 2008go | 23.616 0.075 | 1989B |
| SN 1999ek | 6.639 0.005 | 1989B |
Our agreement (Table 15) with the distances using Cepheids in B by the CCHP is consistent with the findings of Uddin et al. (2024). Those authors obtained H0 72.37 0.71 km s-1 Mpc-1. Concerning TRGB in B, Uddin et al. (2024) gave 70.25 0.71 km s-1 Mpc-1. And for SFB, Uddin et al. (2024) obtained 72.45 0.94.
SN 2001cn, SN 2008go and SN 1999ek give a mean of H0 73.558 2.084 km s-1 Mpc-1.
7.2 Limitations of the parameterization
Pantheon+ obtains the SNe Ia distances in the Hubble flow through a process that implies describing the SNe Ia with SALT2. The CCHP also makes this three rung ladder parameterizing the light curve-rate of decline relation but in a different way.
In this section, we analyse some shortcomings that we have found in the parameterization to obtain the distance moduli through the light curves using either SALT2 or the CSP new light curve-rate of decline relation.
The Pantheon+ light curves are parameterized through the SALT2 light–curve fitting (Tripp 1998; Brout et al. 2022). In this approach the distance modulus is defined as:
where and are global nuisance parameters relating stretch x1 and the color c (respectively) to luminosity. M is the fiducial absolute magnitude of an SN Ia, which can be calibrated by setting an absolute distance scale with primary distance anchors such as Cepheids. The bias is a correction term to account for selection biases that is determined from simulations following Popovic et al. (2021), is the luminosity correction (step) for residual correlations between the standardized brightness of a SN Ia and the host-galaxy mass.
The relation used by the CSP and CCHP instead of being linear in stretch (like in Tripp 1998), is quadratic. The reddening correction is simply a constant, , multiplied by the observed color of the SNe Ia, (B-V) where B and V are the apparent, K-corrected peak magnitudes. P1 is the linear coefficient and P2 is the quadratic coefficient in (sBV-1), which encapsulates the Phillips relation in the sBV stretch computed by SNooPy; Mx would be a fiducial magnitude as in the SALT2 relation in filter x, is the slope of the correlation between peak luminosity and host stellar mass M∗ (see for instance Freedman et al. 2019).
We have noticed that, for reddened SNe Ia, the two parameterizations of the stretch fail to obtain good ’color’ values and this impacts in the distance to SNe Ia. For instance, within the SALT2 parametrization, the color ’c’ in (6) obtained by Pantheon+ tends to be low in the reddened SNe Ia in the Hubble flow (see Figure 14). Such too low ’c’ then reflects in a larger value for the distance modulus and carries into the Hubble flow values of the order of 60 to 65 km s-1 Mpc-1 (Table 15).
This happens as well with the parametrization by the CCHP (see 7), despite the fact that there is a term in (B-V). (B-V) acts as a color term that would include the presence of reddening. In fact, the BV color in the abovementioned CSP parameterization in (7) would be the combination of the intrinsic color of the SN Ia at maximum and the E(B-V) reddening. At maximum the intrinsic (Bmax –Vmax) of normal SNe Ia is already –0.12 mag. However, with the spectra fitted within the ’twin’ comparison one sees that E(B-V) is larger (see Figure 15 and Table 16). In Table 16, we include the E(B-V)MW from Schlafly & Finkbeiner (2011), the derived E(B-V)Tot from the spectra, the ’c’ parameter for this SN Ia in the Pantheon+ compilation, and the ’BV’ parameter in the CCHP compilation.
As mentioned in Freedman et al. (2019), there is an alternative parameterization by Burns et al. (2014) where instead of the (B–V) term there is a RB E(B–V) term to solve better the extinction. However, that has not been implemented in the present derivation of distance moduli for the CSP sample. So, this is the limitation we are seeing.
SN 2007ca and SN 2007A have wrong values of their color either ’c’ or ’BV’ parameters. That causes too large distances and individual points for H0 with low values.
SN 2007A does not have a correct distance modulus in Pantheon+, as it has a color parameter in two diffferent entries of c 0.117 0.024 and c 0.067 0.033. The CSP gives a value of BV 0.201 0.01. The reddening of the spectrum is 0.028 0.012. The CSP gets the right distance modulus in this case as its ’BV’ value would coincide with an E(B-V) from 0.028 0.012 plus a small negative value for the intrinsic color.
SN 2007ca requires a larger color correction than presented in both Pantheon+ and CSP. The color correction in Pantheon+, where it has 3 different entries for the SALT2 parameters is c 0.151 0.023, c 0.204 0.023 and c 0.207 0.028. The E(B-V) for this SN Ia is already 0.3 0.001. The CSP gives as its BV parameter 0.259 0.009, and this time it falls short.
In SN 2001cn, Pantheon+ slightly overcorrects the color with ’c’ 0.118 0.027. The spectra show that E(B-V) is 0.18 0.005. Given that c should reflect the intrinsic color of SNe Ia, which is (Bmax –Vmax) –0.12 mag for normal SNe Ia, if we add the reddening ’c’ it should be around 0.06. Here the final distances are similar, but the one by Pantheon+ has very large errors, that might be linked to not have used the light curves for this SN Ia (Krisciunas et al. 2004) that would allow a very good x1 and c determination.
| SN | ||
| 2012bo | 35.1750.062 | 74.8193.476 |
| 2008bq | 35.8380.058 | 72.052 2.183 |
| 2008bz | 37.235 0.063 | 68.752 2.339 |
| LSQ12fxd | 35.5630.097 | 74.254 3.895 |
| SN 2008bf | 35.100 0.107 | 73.132 4.498 |
| 2007A | 34.310 0.124 | 68.826 5.621 |
| ASASSN-15db | 33.552 0.070 | 67.463 6.475 |
| LSQ14gov | 37.945 0.127 | 73.562 4.137 |
| SN 2007ca | 33.954 0.064 | 73.961 5.113 |
| SN | ||
| 2001cn | 33.9480.066 | 74.2945.421 |
| 2008go | 37.0860.067 | 73.5612.568 |
| 1999ek | 34.3310.069 | 72.9774.743 |
| Galaxy | (CCHP)1 | (SH0ES)2 |
| (mag) | (mag) | |
| M101 | 29.210.03 (stat)0.03 (sys) | 29.120.06 |
| NGC 1365 | 31.380.02 (stat)0.03 (sys) | 31.340.04 |
| NGC 2442 | 31.600.01 (stat)0.04 (sys) | 31.560.05 |
| NGC 4536 | 30.970.01 (stat)0.03 (sys) | 30.930.04 |
| NGC 4639 | 31.730.02 (stat)0.03 (sys) | 31.690.04 |
| NGC 5643 | 30.580.01 (stat)0.04 (sys) | 30.540.04 |
| NGC 7250 | 31.590.02 (stat)0.04 (sys) | 31.590.04 |
| NGC 3972 | 31.670.04 (stat)0.03 (sys) | — |
| NGC 4038 | 31.530.06 (stat)0.04 (sys) | — |
| NGC 4424 | 31.150.04 (stat)0.03 (sys) | — |
| 1Lee et al. (2025). | 2Li et al. (2025) |
7.3 The value of H0
Our results for individual SNe Ia, in Tables 15 and 18 clearly confirm that the Hubble tension is real, since they are incompatible with the H0 value derived from the CMB. The value, for the 12 SNe Ia, is H0 = 72.833 1.306 (stat) 1.151 (syst) km s-1 Mpc-1 when calibrating the anchor distances with Cepheids (by SH0ES) and 72.388 1.272 (stat) 1.015 (sys) km s-1 Mpc-1 when calibrating with JAGB (by the CCHP). The average is H0 = 72.610 1.289 (stat) 1.085 (sys) km s-1 Mpc-1, far from the 67.4 0.5 km s-1 Mpc-1 of the Planck Collaboration.
Figure 16 is the Hubble diagram of the SNe Ia in this work anchored in Cepheids (Table 15). The slope corresponds to H0 72.833 km s-1 Mpc-1. The slope of the Hubble diagram, when the SNe Ia are anchored in JAGBs (CCHP) (Table 18), is quite similar with H0 of 72.388 km s-1 Mpc-1.
If we take the SNe Ia of the BL class only, for which the anchor distance is known with the greatest precision (coincidence between the Cepheid distances from Freedman et al. (2019) and Saha et al. (1999), the value grows to H0 = 73.556 2.084 (stat) km s-1 Mpc-1 (weighted average from the three SN Ia), still farther from the Planck value.
Our results from comparing the spectra of very nearby SNe Ia with those of their twins in the Hubble flow also validates the three rung procedures used by the Pantheon+ and CCHP collaborations, even at the level of individual SNe Ia, since the discrepancies can easily be explained by differences in the evaluation of the total reddenings or the different light-curve parameters adopted, as we have seen in the previous Section.
To leave our results stable against future changes in the distances of the anchors in the nearby galaxies, we publish the distance factors between the anchors and the SNe Ia distances. The factors should be multiplied by the distances of the anchors to obtain the distances to the SNe Ia in the Hubble flow (Table 17).
Concerning the different calibrations used for setting our anchor distances, we see that the distance moduli from TRGB by the CCHP have fluctuated during the last year. In addition, their distances are very different when using the same data but analysed by other groups (Anand et al. 2022).
Consequently, we have obtained the two H0 values given above, for the two different calibrations of our anchors using Cepheids and JAGBs (Table 15 and Table 18).
It is worth to remark the distance to M101 measured from SH0ES with Cepheids and with the JAGB stars from both the CCHP and SH0ES. Those distance values clearly favor a value below 70 km s-1 Mpc -1 for H0 from SNe Ia twins of SN 2011fe. Table 19 illustrates the difference in distance moduli to M101 obtained by the two collaborations with the same method: it is of 0.09 mag, which translates in 4 km s-1 Mpc-1 for H0. If our sample would be dominated by 2011fe–like SNe Ia, the total H0 would be lower than what it is. However, the other values of distances using JAGB stars for the rest of galaxies differ only by 0.04 mag between the larger distance value by CCHP and the shorter one from SH0ES. As the sample contains SNe Ia twins from SN 2013aa and SN 2013dy, the values above 70 km s-1 Mpc -1 dominate. We pointed, in RLGH24, a preference for the distance to SN 2011fe from the spectra of 6.5 Mpc, in agreement with the value obtained from Blue Supergiants (Bresolin et al. 2022) of 6.5 0.2 Mpc. Inevitable, using the 10 galaxy calibrators by the JAGB in rung 2, as it is done by the CCHP, could bring H0 to lower values than 70 km s-1 Mpc-1 given the calibration of M101. That would depend on the weight of M101 in the whole sample. However, it is surprising such low mean value of H0 67 km s-1 Mpc-1 published by Freedman et al. (2025). The case deserves further exploration.
8 Summary and conclusions
The use of the ”SNe Ia twins” to obtain H0 has revealed that the Hubble tension is real. The distances obtained through this direct path go from 6 Mpc to around 300 Mpc in a single step.
Comparing our distances with those in the Pantheon+ and CCHP samples using Cepheids, we find a general agreement, with a few exceptions, mostly related with the determination of the reddening of some SNe Ia. The slopes in the Hubble diagram coincide and the values of H0 are very close.
This implies that the SNe Ia calibrators in rung 2 are a good sample that achieves a consistent final value of H0 in rung 3. The three–rung method seems to work. It could fail if a biased sample in rung 2 were taken. Then, the final value of H0 would be distorted, and the slope of the Hubble diagram changed. The large samples in Pantheon+ or CCHP using Cepheids have values of individual H0 whose average gives close values for H0. Using the TRGB obtained by the CCHP, the distances of the galaxies in rung 2 would be larger and the H0 smaller, in contrast to the TRGB from SH0ES. A measurement that uses Cepheids from SH0ES and JAGB from the CCHP with the twins method crosses the aisle between the two collaborations. The opportunity comes from similar distance values for a few galaxies below 20 Mpc.
At the individual level, the distances to each SN Ia are affected by the limitations of the parameterization of the light curves used by the two collaborations. The method of twin SNe Ia, with its use of spectra, is very powerful in getting the reddening of the SNe Ia in the Hubble flow. An approach as that developed here can point to the failures in the distance moduli for these supernovae that give H0 in the extremes of the distribution (very low or very large H0). SNe Ia with color larger than ’c’ 0.1 in the Pantheon+ sample give rise to low reddening correction and larger distance moduli. This leads to individual points of H0 falling out of the range of discussion, i.e., in the tail of H0 below 65 km s-1 Mpc-1, A similar effect is seen with the BV color. Those tails come from SNe Ia where the color terms ’c’, or ’(BV)’ are not well estimated or even the rates of decline of the lightcurves parameterized in the form of ’x1’ in SALT2 or sBV in CCHP have been wrongly estimated. Sometimes large uncertainties are quoted to accompany those values, but in some cases they are not. Those extreme values do not alter significantly the mean for H0. However, for those particular SNe Ia in the Hubble flow the distance remains unknown.
Inaccurate point values in distances and therefore in H0 individual points (even when enclosing these points within a a large error bar) would interfere with the proper determination of possible cosmic bulk flow in the z range 0.015 z 0.1. An important task is to improve the Pantheon+ and CSP distances. In this sense, our approach can help a lot to give the right distances and right individual H0 values.
The use of SN 1989B in M66 as anchor has been very fruitful. It has allowed to get accurate distances to three SNe Ia: SN 2001cn, SN 2008go and SN 1999ek giving rise to a H0 of 73.556 2.084 km s-1 Mpc-1 (weighted average of the consistent three values). As said before, the distance to M66 with TRGB by Freedman et al. (2019) coincides with the Cepheids value from Saha et al. (1999).
We finally stress that the present use of two different calibrations, that with Cepheids from Pantheon+ and the one with JAGB from CCHP, do converge to the same result. This concordance found using anchors from the two collaborations rarely happens and looks very promising. There is a very good agreement on the distances to NGC 7250 and NGC 5643 derived with Cepheids by SH0ES and derived with the use of J-Asymptotic Giant Branch stars (JAGB stars) by the CCHP. The difference for SNe Ia in the Hubble flow anchored in one or the other way is negligible: our SNe Ia sample in the Hubble flow anchored in Cepheids gives H0 72.833 1.306(stat) 1.151 (sys) and anchored in JAGB stars gives H0 72.388 1.272 (stat) 1.015 (sys). The mean would be H0 72.610 1.289 (stat) 1.085 (sys).
This work has made use of spectra from the Weizmann Interactive Supernova data REPository (WISeREP), spectra and photometry of the Carnegie Supernova Project (CSP). The authors acknowledge Pantheon+ and the CSP for opening their data releases and software to the community. We are very grateful to Nidia Morrell (from the CSP) for exchanges in relation to the photometry of CSP SNe Ia. PR-L and A Q-E acknowledge support from grant PID2021-123528NB-I00, from the the Spanish Ministry of Science and Innovation (MICINN). JIGH acknowledges financial support from MICIU grant PID2023-149982NB-I00. A.P. acknowledges support from the PRIN-INAF 2022 project “Shedding light on the nature of gap transients: from the observations to the model”. .
Software: AstroPy (Price-Whelan et al. 2018); H0CSP.ipynb (Uddin et al. 2024); Matplotlib (version 3.2.1, Hunter 2007); NumPy (version 1.18.2, Oliphant 2006; van der Walt et al. 2011); SNooPy (Burns et al.2011).